OpenStudy (anonymous):

help to solve 61-2{3[12-(13-4)/3]}

OpenStudy (kc_kennylau):

Calculate the stuff inside (parentheses) first, then the stuff inside [brackets]

OpenStudy (kc_kennylau):

61-2{3[12-(13-4)/3]} =61-2[3(12-9/3)] =61-2[3(12-3)] =61-2[3(9)] =61-2(27) =61-54 =9

OpenStudy (kc_kennylau):

PEMDAS: P - parentheses E - exponents MD - multiplication and division; whichever comes first AS - addition and subtraction; whichever comes first